OpenStudy (anonymous):

simplify 2^3X2^5=2^15 right

OpenStudy (anonymous):

You add the exponents. 2^8

OpenStudy (anonymous):

The reason is you have \[(2x2x2)x(2x2x2x2x2)=2^8\]

OpenStudy (anonymous):

i dont get it:(

OpenStudy (anonymous):

\[2^3\times2^5=(2\times2\times2)\times(2\times2\times2\times2\times2)=2\times2\times2\times2\times2\times2\times2\times2=2^8\]

OpenStudy (anonymous):

The first set of parentheses is the expansion of 2^3, and the second set is the expansion of 2^5. If you take away the parentheses, you get eight 2s multiplied together.